OpenStudy (countrygirl1431):

Solve the following equation. Then place the correct number in the box provided. 2(x + 1 ) - 3(x + 5) ≥ 0

OpenStudy (trisaba):

\[2\left( x +1 \right)=2\left( x \right)+2\left( 1 \right)\]

OpenStudy (trisaba):

\[-3\left( x +5 \right)=-3\left( x \right)+-3\left( 5 \right)\]

OpenStudy (anonymous):